Find the length of X. Express in simplified radical form.
I will try to draw this out so it makes it easier:
-
- -
- -
- -
51ft - - X
- -
- -----
- - -
----------------------------------
45 ft
How do I go about solving this type of equation? Thank you so very much!

Answer by ankor@dixie-net.com(12692)

Find the length of X. Express in simplified radical form.
I will try to draw this out so it makes it easier:
-
- -
- -
- -
51ft - - X
- -
- -----
- - -
----------------------------------
45 ft
How do I go about solving this type of equation? Thank you so very much!
:
I assume this is a right triangle, one side is 45 ft, hypotenuse is 51,
The problem is to find the other side (x)
:
You should know the Pythagorean Theorem: a^2 + b^2 = c^2
:
In your problem a = 45 and c = 51, b = x so we have this equation:
45^2 + x^2 = 51^2
:
2025 + x^2 = 2601; squared the numbers
:
x^2 = 2601 - 2035; subtract 2025 from both sides
:
x^2 = 576
:
x = Sqrt(576)
:
x = 24 ft
:
:
Check: 45^2 + 24^2 = 2601 which is 51^2